OpenStudy (anonymous):

7x+7-9x=14-15x+13x-7 A. the solution to x = ? B. The solution is all real numbers C. There is no solution

OpenStudy (anonymous):

7x+7-9x=14-15x+13x-7 (7x-9x+15x-13x) = (14-7-7) add together the coefficients and post what you get.